There are 4 red, 3 green and five blue discs.two discs are drawn out. Find probability that two discs of same color are drawn out

1
Expert's answer
2021-12-27T10:54:38-0500

There are 4 red, 3 green and five blue discs.

Total number of discs = 4 + 3 + 5 = 12

Number of ways of drawing 2 discs from 12 discs = "12\raisebox{0.25em}{$C$}2"

If two discs are drawn out then the probability that two discs of same color are drawn out is:

"P=\frac{(4\raisebox{0.25em}{$C$}2)+(3\raisebox{0.25em}{$C$}2)+(5\raisebox{0.25em}{$C$}2)}{12\raisebox{0.25em}{$C$}2}"


"P=\frac{2.92}{66}=0.044"


There is 4.4 percent probability that two discs of same color are drawn out from 12 discs.